OpenStudy (anonymous):

What is the derivative of (x^2+1)^x ?

OpenStudy (abb0t):

I suggest logrithmig differentiation.

OpenStudy (abb0t):

\[y = (x^2+1)^x \] take the derivative of that bad boy! all of it. even y.

hartnn (hartnn):

like 1st step will be \(\log y =\log(x^2+1)^x \\ \log y=x \log(x^2+1)\) now there is use of product rule and chain rule on right side....